Delilah opens a credit card with an APR of 24.38% compounded monthly. How much is charged in interest this month if her balance is $550?
Answer by jim_thompson5910(35256)

24.38% compounded monthly ---> 24.38/12 = 2.031667% monthly interest rate

2.031667% of 550 = 0.02031667*550 = 11.1741685 = 11.17

The amount charged in interest is roughly $11.17